OpenStudy (anonymous):

Algebra 2 Connexus students!!!! I WILL DO ANYTHING... PLEASE HELP

OpenStudy (anonymous):

In your lab, a substance's temperature has been observed to follow the function T(x)=(x+5)^3, The turning point of the graph is where the substance changes from a liquid to a solid. Explain to your fellow scientists how to find the turning point of this function, USING COMPLETE SENTENCES.

OpenStudy (anonymous):

Well first of all, you need to factor that equation into a polynomial. So factor: (x+5)^3 (x+5)(x+5)(x+5) x^3+15x^2+75x+125

OpenStudy (anonymous):

But notice in the part of factoring where it says (x+5)(x+5)(x+5), there is x+5 three times.

OpenStudy (anonymous):

There for there is only one repeating zero of the polynomial, and that zero is "encoded" in x+5. So by setting x+5=0, you turn out with -5 as the zero.

OpenStudy (anonymous):

i have to write it complete sentences.. btw Thank you for helping me!

OpenStudy (anonymous):

So this polynomial turns at -5

OpenStudy (anonymous):

@evan97 do you mind helping me some more?

OpenStudy (anonymous):

Sure

OpenStudy (anonymous):

two students in your class, tucker and karly are disputing a function. tucker says that for the function, between x=-3 and x=3, the average rate of change is 0. Karly says that for the function, between x=-3 and x=3, the graph goes up through a turning point, and then back down. Explain how tucker and karly can both be correct, USING COMPLETE SENTENCES

OpenStudy (anonymous):

Graphing the function?

OpenStudy (anonymous):

Would that be acceptable?

OpenStudy (anonymous):

I have to use sentences

OpenStudy (anonymous):

They can graph the function using paper or a calculator to determine where the turning point lies.

OpenStudy (anonymous):

can you expand more on how they do that? :)

OpenStudy (anonymous):

@evan97

OpenStudy (anonymous):

but if thats all.... can we move to another question? :)

OpenStudy (anonymous):

Yeah sure. Explaining how to graph polynomials is a long process so you might just want to leave it at that. Next question?

OpenStudy (anonymous):

you are having a conference call with the CEO of a paper company. You have interpreted the number of trees cut down versus profit as the function P(x)=-x^4+x^3+7x^@-x. describe to the CEO what the graph looks like. Use COMPLETE SENTENCES and focus on the end behaviors of the graph AND WHERE THE COMPANY WILL BREAK EVEN.

OpenStudy (anonymous):

The part in caps(the second part of the question) is what i need help with

OpenStudy (anonymous):

@evan97

OpenStudy (anonymous):

hello @evan97 ??

OpenStudy (anonymous):

I hate flvs .. lol

OpenStudy (anonymous):

but im in it so xD

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

anybody red badge of courage help please @nicole96 @even97 @STUCKINTHIS

OpenStudy (anonymous):

@evan97